OpenStudy (anonymous):

4. What is the slope of a line parallel to the line with equation 5x + 3y = 7? (Points : 4)

OpenStudy (anonymous):

Can you put that equation into the form y = mx + b? Just try and isolate that y on the left side of the equation.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

equation 5x + 3y = 7

jimthompson5910 (jim_thompson5910):

Similar Example 2x + 3y = 10 Solve for y to get \[\large 2x + 3y = 10\] \[\large 3y = -2x+10\] \[\large y = -\frac{2}{3}x+\frac{10}{3}\] So the slope here is -2/3 Note: this is NOT the answer as it's just an example

OpenStudy (anonymous):

thanks..

jimthompson5910 (jim_thompson5910):

so using that example, what's the slope of the line 5x + 3y = 7 ?

OpenStudy (anonymous):

Or for other standard equations Ax=By=C slope is -b/a so slope of 5x+3y=7 would be -3/5

OpenStudy (anonymous):

whaaa. ugh i hate math.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

ya i do i get stuck on ths part hold up imma draw real quick...

jimthompson5910 (jim_thompson5910):

jdifatta, the slope of 5x + 3y = 7 is NOT -3/5

OpenStudy (anonymous):